# TopoStreamer: Temporal Lane Segment Topology Reasoning in Autonomous Driving

### v1

@@ Abstract
Lane segment topology reasoning constructs a comprehensive road network by capturing the topological relationships between lane segments and their semantic types. This enables end-to-end autonomous driving systems to perform road-dependent maneuvers such as turning and lane changing. However, the limitations in consistent positional embedding and temporal multiple attribute learning in existing methods hinder accurate roadnet reconstruction. To address these issues, we propose TopoStreamer, an end-to-end temporal perception model for lane segment topology reasoning. Specifically, TopoStreamer introduces three key improvements: streaming attribute constraints, dynamic lane boundary positional encoding, and lane segment denoising. The streaming attribute constraints enforce temporal consistency in both centerline and boundary coordinates, along with their classifications. Meanwhile, dynamic lane boundary positional encoding enhances the learning of up-to-date positional information within queries, while lane segment denoising helps capture diverse lane segment patterns, ultimately improving model performance. Additionally, we assess the accuracy of existing models using a lane boundary classification metric, which serves as a crucial measure for lane-changing scenarios in autonomous driving. On the OpenLane-V2 dataset, TopoStreamer demonstrates significant improvements over state-of-the-art methods, achieving substantial performance gains of +3.4% mAP in lane segment perception and +2.1% OLS in centerline perception tasks. Our code will be released.
@@ Introduction
Perception serves as a crucial component in end-to-end autonomous driving (Li et al. 2024b; Yang et al. 2025), providing essential road priors for planning. Existing HD map learning and lane topology reasoning methods primarily focus on frame-by-frame detection (Li et al. 2023b; Liao et al. 2022). This approach fails to account for instance consistency across consecutive frames, making it susceptible to missed detections due to occlusions and high-speed movements (Yuan et al. 2024). Such limitations significantly hinder continuous and smooth decision-making and maneuvers. To comprehensively leverage temporal information, streaming-based methods (Yuan et al. 2024; Wang et al. 2024b; Wu, Yang, and Li 2025) propose memory-based temporal propagation to establish long-term frame associations. Specifically, these approaches leverage the ego-vehicle pose to predict the probable positions of road instances in subsequent frames. However, they fail to capture sufficient road information for planning. This inspired us to introduce a temporal mechanism in lane topology reasoning, which we can leverage perception and topology reasoning results from previous frames to predict current frame outcomes and capture fine-grained road information for planning (Jia et al. 2025). Fig. 1 demonstrates the comparison between our method and current streaming-based learning methods(Yuan et al. 2024). To the best of our knowledge, achieving this objective presents two primary challenges: (1) Consistent positional embedding. Current streaming-based methods exhibit deficiencies in their positional embedding design for stream queries. Furthermore, certain lane topology reasoning approaches (Li et al. 2023a, b) suffer from inconsistency between reference points and positional embedding updates. (2) Temporal multiple attribute learning for lane segments. Topology reasoning between lanes is highly sensitive to the precise localization of lane segments (Fu et al. 2025a) and projection errors make it challenging to maintain consistent localization and category of lane segments across temporal propagation.
@@
TopoStreamer (ours) 24 Yes 36.6 $\uparrow 3.4\%$ 35.0 $\uparrow 2.0\%$ 38.1 $\uparrow 4.7\%$ 28.5 $\downarrow 2.3\%$ 50.0 $\uparrow 4.1\%$ 13.6
@@
TopoStreamer (ours) 24 Yes 44.4 $\uparrow 2.1\%$ 35.2 $\uparrow 0.5\%$ 28.8 $\uparrow 4.1\%$
Table 2: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on centerline perception. * denotes metrics from the $\text{TOP}_{ll}$ v 1.0.0 version extracted from the referenced paper. TopoFormer utilizes a pretrained lane detector.
@@
Results on Lane Segment The results for lane segment are shown in Tab. 1. Compared with single-frame detection methods, we outperforms LaneSegNet by 4.0% mAP, 3.1% TOPlsls and 4.1% Accb, and exceeds TopoLogic by 3.4% mAP. This shows the effectiveness of our streaming design for lane segment. Compared with temporal detection methods, we achieve a remarkable improvement of 10.6% mAP, 11.9% TOPlsls and 10.6% Accb. They exhibit limitations in detecting more fragmented lane segments, as they fail to account for multiple attributes and PE design.

### v2

@@ Abstract
Lane segment topology reasoning constructs a comprehensive road network by capturing the topological relationships between lane segments and their semantic types. This enables end-to-end autonomous driving systems to perform road-dependent maneuvers such as turning and lane changing. However, the limitations in consistent positional embedding and temporal multiple attribute learning in existing methods hinder accurate roadnet reconstruction. To address these issues, we propose TopoStreamer, an end-to-end temporal perception model for lane segment topology reasoning. Specifically, TopoStreamer introduces three key improvements: streaming attribute constraints, dynamic lane boundary positional encoding, and lane segment denoising. The streaming attribute constraints enforce temporal consistency in both centerline and boundary coordinates, along with their classifications. Meanwhile, dynamic lane boundary positional encoding enhances the learning of up-to-date positional information within queries, while lane segment denoising helps capture diverse lane segment patterns, ultimately improving model performance. Additionally, we assess the accuracy of existing models using a lane boundary classification metric, which serves as a crucial measure for lane-changing scenarios in autonomous driving. On the OpenLane-V2 dataset, TopoStreamer demonstrates significant improvements over state-of-the-art methods, achieving substantial performance gains of +3.0% mAP in lane segment perception and +1.7% OLS in centerline perception tasks.
@@ Introduction
Perception serves as a crucial component in end-to-end autonomous driving (Li et al. 2024b; Yang et al. 2025b), providing essential road priors for planning. Existing HD map learning and lane topology reasoning methods primarily focus on frame-by-frame detection (Li et al. 2023b; Liao et al. 2022). This approach fails to account for instance consistency across consecutive frames, making it susceptible to missed detections due to occlusions and high-speed movements (Yuan et al. 2024). Such limitations significantly hinder continuous and smooth decision-making and maneuvers. To comprehensively leverage temporal information, streaming-based methods (Yuan et al. 2024; Wang et al. 2024b; Wu, Yang, and Li 2025) propose memory-based temporal propagation to establish long-term frame associations. Specifically, these approaches leverage the ego-vehicle pose to predict the probable positions of road instances in subsequent frames. However, they fail to capture sufficient road information for planning. This inspired us to introduce a temporal mechanism in lane topology reasoning, which we can leverage perception and topology reasoning results from previous frames to predict current frame outcomes and capture fine-grained road information for planning (Jia et al. 2025). Fig. 1 demonstrates the comparison between our method and current streaming-based learning methods(Yuan et al. 2024). To the best of our knowledge, achieving this objective presents two primary challenges: (1) Consistent positional embedding. Current streaming-based methods exhibit deficiencies in their positional embedding design for stream queries. Furthermore, certain lane topology reasoning approaches (Li et al. 2023a, b) suffer from inconsistency between reference points and positional embedding updates. (2) Temporal multiple attribute learning for lane segments. Topology reasoning between lanes is highly sensitive to the precise localization of lane segments (Fu et al. 2025a) and projection errors make it challenging to maintain consistent localization and category of lane segments across temporal propagation.
@@
TopoStreamer (ours) 24 Yes 36.6 $\uparrow 3.0\%$ 35.0 $\uparrow 1.3\%$ 38.1 $\uparrow 4.6\%$ 28.5 $\downarrow 2.3\%$ 50.0 $\uparrow 1.9\%$ 13.6
@@
TopoStreamer (ours) 24 Yes 44.4 $\uparrow 1.7\%$ 35.2 $\uparrow 0.5\%$ 28.8 $\uparrow 1.8\%$
Table 2: Comparison with the state-of-the-arts on OpenLane-V2 benchmark on centerline perception. TopoFormer adopts a staged training strategy that utilizes a pretrained lane detector for topology reasoning training. While this leads to better detection performance, it offers only slight advantage in topology prediction.
@@
Results on Lane Segment The results for lane segment are shown in Tab. 1. Compared with single-frame detection methods, we outperforms LaneSegNet by 4.0% mAP, 3.1% TOPlsls and 4.1% Accb, and exceeds Topo2Seq by 3.0% mAP. This shows the effectiveness of our streaming design for lane segment. Compared with temporal detection methods, we achieve a remarkable improvement of 10.6% mAP, 11.9% TOPlsls and 10.6% Accb. They exhibit limitations in detecting more fragmented lane segments, as they fail to account for multiple attributes and PE design.